OpenStudy (anonymous):

Roots of quadratic:

OpenStudy (anonymous):

if α and β are the roots of equation 4x^2 - 2x - 1 = 0, find α - β

OpenStudy (raden):

use formula : α - β = | (sqrt(D)/a) | with D=b^2-4ac

OpenStudy (anonymous):

oh yeah... forgot to put - without finding the roots So like using αβ = c/a and α+β=-b/a

OpenStudy (raden):

right, but ur problem to find value of α - β?

OpenStudy (anonymous):

yes. find α - β

OpenStudy (raden):

we know that general form of a quadratic equation is ax^2+bx+c=0 so, we can find α - β with formula : α - β = abs(sqrt(D)/a)

Parth (parthkohli):

You can always count on finding the roots and assuming them as \(\alpha, \beta\).

Parth (parthkohli):

\[x = {2 \pm \sqrt{4 - 4(4)(-1)} \over 8}\]

OpenStudy (anonymous):

@ParthKohli, without finding roots

OpenStudy (raden):

yea, right to.. but should use absolut sign because not known α > β or α < β

OpenStudy (anonymous):

yeah ok

OpenStudy (raden):

so, what ur answer ?

OpenStudy (anonymous):

1.128

OpenStudy (raden):

yea, or (sqrt(5))/2 ?